OpenStudy (anonymous):

Describe the steps you would take to write and equation in point slope form of the line that passes through the points (3.-2) and (4,5)

OpenStudy (saifoo.khan):

Begin with finding the slope.

hero (hero):

Evil Double Zero, help him

OpenStudy (saifoo.khan):

Max, do u know how to find the slope?

OpenStudy (anonymous):

ya y2-y1/x2-x1

OpenStudy (saifoo.khan):

Right, find it.

OpenStudy (anonymous):

7/1 7

OpenStudy (anonymous):

slop=7

OpenStudy (saifoo.khan):

Right.

OpenStudy (saifoo.khan):

Now, my way of solving is.. use one point and a slope, insert the values in y=mx+b this will help you get the y-intercpt or 'b'.

OpenStudy (anonymous):

no i figured it out in point slope form it is y-2=7(x-3)

OpenStudy (anonymous):

y+2 actually sorry

OpenStudy (saifoo.khan):

Great.

OpenStudy (saifoo.khan):

that's the answer.